OpenStudy (anonymous):

PLEASE help with solving systems of linear equations with three variables! x + 3y + 2z = -2 2x - y - z = -9 x- 2y + 5z = 1

OpenStudy (anonymous):

this one has a long process. It's similar to that of system of equations with two variables, just you have to repeat the process more times. You solve for one variable, say x in the first equation. that equation in to the second one, leaving you with y and z. Now solve for y, and plug it in to the third one, leaving you with just z. You should get a solid number now, and you can plug in that value into the first problem. Basically repeat that until you have all three values.

OpenStudy (thomas5267):

Are you studying linear algebra or just algebra in high school?

OpenStudy (anonymous):

Do you need help???

OpenStudy (anonymous):

Yes , I still do not understand. Whenever I try this problem i get getting huge fractions like x= 150/52

OpenStudy (anonymous):

ok, do it like me. For example: from first equation: x = -2 -3y-2z -> from second: 2(-2 -3y-2z) -y -z = -y

OpenStudy (thomas5267):

The answer is neat integers. Try again.

OpenStudy (anonymous):

from second equation you must find y or z and than put it to third equation

OpenStudy (anonymous):

x = -2 -3y-2z -> from second: 2(-2 -3y-2z) -y -z = -9

OpenStudy (radar):

Yes, the solutions are small integers, with one variable actually being a "0"